OpenStudy (anonymous):

if we say z1 / z2: z1 - 2√2 ( cos(2π/3)+i sin2π/3)) z2- 2√3 ( cos (π/6)+i sin(π/6)) is the answer √6 / 3 or √6 /3 I?

OpenStudy (zehanz):

If you mean: \(z_1=2\sqrt{2}(\cos \frac{2\pi}{3}+i\sin\frac{2\pi}{3})\) and \(z_2=2\sqrt{3}(\cos \frac{\pi}{6}+i\sin\frac{\pi}{6})\), then \(\dfrac{z_1}{z_2}=\dfrac{2\sqrt{2}(\cos \frac{2\pi}{3}+i\sin\frac{2\pi}{3})}{2\sqrt{3}(\cos \frac{\pi}{6}+i\sin\frac{\pi}{6})}=\dfrac{\sqrt{6}(\cos \frac{2\pi}{3}+i\sin\frac{2\pi}{3})}{3(\cos \frac{\pi}{6}+i\sin\frac{\pi}{6})}\). Now before doing anything algebraic with this number, remember what dividing two complex numbers means geometrically: you have to divide their magnitudes (absolute values, moduli, or whatever you call them...) and subtract their arguments. We have already divided the moduli: \(|\dfrac{z_1}{z_2}|=\dfrac{\sqrt{6}}{3}\). Now we have to subtract the arguments. We see this: \(\arg(z_1)=\frac{2\pi}{3}=\frac{4\pi}{6}\), \(\arg(z_2)=\frac{\pi}{6}\). So all you have to do is: \(\arg\left( \dfrac{z_1}{z_2}\right)=\frac{4\pi}{6}-\frac{\pi}{6}=...\) As you can see, the number you are looking for is very simple indeed! Can you see now what it is?

OpenStudy (anonymous):

the answer is then 3π / 6?

OpenStudy (zehanz):

The argument of \(\dfrac{z_1}{z_2}\) is 3π/6 = π/2. That puts is on the imaginary axis. (π/2 = 90 degrees). Because we've found the magnitude to be \(\frac{\sqrt{6}}{3}\), the answer must be: \(\frac{\sqrt{6}}{3}i\)

OpenStudy (anonymous):

ok so it must have a I !! thanks

OpenStudy (zehanz):

YW!